NATIONAL INCOME ACCOUNTS Use the data below to answer questions 1-3: Consumption Expenditures Wages Rent Interest Income Corporate Profits Proprietors Income Farm Income Personal Income Personal Taxes $8,000 billion $7,000 billion 150 billion S 500 billion S 500 billion $800 billion S 150 billion $9,000 billion $ 900 billion 1. National income (NI) is equal to $7,000 billion $8,200 billion $9,900 billion $9,100 billion b. d. 2. Disposable income (DI) is equal to $7,100 billion $8,100 billion $9,100 billion $9,900 billion a. b. d. Personal saving is equal to (3. a. $1,000 billion $900 billion $100 billion $50 billion C. d. a. that is taken out of your pay and disposed of by government. b. available to households for spending and saving. c. earned by the factors of production in a given time period. 4. Disposable income is the income: d. received by persons from all sources before payment of person only need help on 3

If, in a given year, nominal GDP is $11,850 billion and the GDP deflator is 120, then real GDP is: a. $8,500 billion. b. $14,220 billion. c. $9,875 billion. d. $9,500 billion.e expenditure approach to calculating GDP is represented by the e quation: GDP -C+I+ G- IM. GDP = C + I + G + (EX-IM) GDP NI+ net foreign factor income. GDP = C + I + G + (EX + IM). oss Domestic Product measures the: nations well-being. total market value of all final goods and services produced within the U total market value of all final goods and services produced by U.S. cit market value of all intermediate goods plus the market value of all final goods produced in the U.S. .S izens.
Answer #1

3) Personal Saving = Disposable Income - Consumption Expenditure = $ (8100 - 8000) = $100

Option B is correct.

GDP is the total market value of all the final goods and services produced within a country in a given period of time.

According to Expenditure Method of National Income,

GDP = C + I + G + ( X - M)

G D Pole f lator = (Ν omínalG D P / Real G D P) * 100

RealGDP (11850 / 120)*100 = 9875

Option C is correct.

Disposable Income is the income that is available to the household for spending and saving because this income is left with the individual after paying all the taxes.